# Distributed Massive MIMO System with Dynamic Clustering in LEO Satellite Networks

### v1

@@ Abstract
Low Earth Orbit (LEO) satellite networks, distributed mMIMO, User-centric, asynchronous transmission.
@@ I Introduction
Multiple input multiple output (MIMO) is a technology implemented to enhance the performance in wireless communication systems [1]. However, co-located MIMO in satellites communication fails to provide the same performance as that in terrestrial networks [2]. This is because satellite channels generally differ from those of the terrestrial network. In particular, strong line-of-sight (LoS) channel is one of the main characteristics of satellite channels due to the lack of scatters in the space. This satellite channel characteristic results in high spatial correlation when co-located MIMO systems are used, making it difficult for the receiver to estimate the channel coefficient for each antenna and hence affecting the validity of the channel state information which is necessary for MIMO system operation. As such, satellite networks, have the drawback in that they are unable to effectively exploit the performance enhancement provided by MIMO technology compared to terrestrial networks [3]. To overcome this issue, instead of using co-located MIMO, a distributed MIMO system is proposed which is able to provide uncorrelated channels among spatially separated antennas [4].
Distributed MIMO in space requires satellite cooperation to implement beamforming or joint transmission and reception. In [5], the authors proposed a joint transmission with two LEO satellites where it is proved that the relative location between the cooperative satellites and the ground user significantly affects the system performance due the fact that LOS channel is dominant. The problem of overlap in satellite coverage areas is studied in [6] where a user receives multiple signals in a multi-satellite scenario. In [4] a distributed massive MIMO (mMIMO) system is proposed for LEO satellite networks where multiple satellite access points cooperate to serve a set of ground users. For the backhauling system, a central super satellite with advanced computing capabilities is used as a central unit in order to coordinate resources and control satellite cooperation. A theoretical framework to investigate the spectral efficiency performance of distributed mMIMO in space was developed in [7]. The results in [4] and [7] show that in addition to improving the average system spectral efficiency, distributed mMIMO can reduce the handover rate since the user receives its data from a cluster of serving satellites instead of a single satellite connection. In [8], a distributed beamforming technique was proposed to support direct communications from LEO satellites to smartphones.
@@
In this paper, we propose a scalable distributed mMIMO system designed for LEO satellite networks, utilizing dynamic user-centric clustering. The proposed system is called a user-centric distributed mMIMO (UC-DMIMO) system. With a focus on achieving scalable implementation, we provide new algorithms for initial cooperative access, pilot assignment, and cluster selection. To further enhance system efficiency in terms of complexity and handover rate, each user is allocated a reference satellite access point (RSAP) responsible for exchanging the user’s information with other satellites in its designated serving cluster (SC). Two distinct approaches are provided for RSAP selection: either selecting the LEO satellite with the best channel or choosing the one with the maximum service time. Furthermore, phase-aware pre-coding is used to synchronize the phases of signals received from different spatially distributed antennas. Finally, the system performance in terms of spectral efficiency and the SC size of the proposed UC-DMIMO system is compared with that of the baseline full-cooperative DMIMO (FC-DMIMO) systems, where all satellites cooperatively serve all users simultaneously, as well as non-cooperative transmission (NCT) systems, where each user is connected to only a single satellite.
@@ II-B Channel Model
$\textbf{h}_{nm}=\beta_{nm}\textbf{g}_{nm}$ (1)
@@
where $\kappa$ refers to the Rician K-factor and $\textbf{g}^{\prime}_{nm}$ and $\textbf{g}^{\prime\prime}_{nm}$ corresponds to the LOS and NLOS components. Specifically, $\textbf{g}^{\prime\prime}_{nm}$ characterizes the random component, accounting for the scattered multipath that follows a Rayleigh distribution its elements are independent and identically distributed random variables with a zero mean and unit variance. Besides, according to [10], the $l$-th element of $\textbf{g}^{\prime}_{nm}$ is given as
@@
$\beta_{nm}=\frac{1}{{\mathcal{L}^{prop}_{nm}}\mathcal{L}^{sh}_{nm}\mathcal{L}^{ant}_{nm}\mathcal{L}^{other}_{nm}}$ (4)
@@
${\mathcal{L}^{prop}_{nm}}[\textrm{dB}]=32.45+20\textrm{log}_{10}f+20\textrm{log}_{10}r_{nm}$ (5)
The parameters $\mathcal{L}^{sh}_{nm}$ in (4) represents the losses due to the shadowing effect, which is modeled as a log-normal random variable, i.e., $\mathcal{L}^{sh}_{nm}\sim\mathcal{N}(0,\sigma_{sh}^{2})$, where $\sigma_{sh}^{2}$ is the shadowing variance. $\mathcal{L}^{ant}_{ml}$ denotes the antenna gain losses due to beam misalignment . Based on the 3GPP report in [9], this losses is modeled as
$\mathcal{L}^{ant}_{nm}=\frac{1}{4}\bigg{|}\frac{2\pi\eta\sin{w_{nm}}}{J_{1}(2\pi\eta\sin{w_{nm}})}\bigg{|}^{2}$ (6)
@@ II-C Asynchronous Reception
$\theta_{nm}=\exp{\big{(}-2\pi\frac{\Delta t_{nm}}{T_{s}}\big{)}}$ (7)
@@ III-B Channel Estimation
$\mathbb{Y}^{P}_{m}=\sum_{n=1}^{N}\sqrt{a_{nm}p_{n}}\tilde{\mathbb{h}}_{nm}\mathbb{\Xi}_{n}^{T}+\mathbb{W}_{m}$ (9)
@@
$\!\!\!\!\!\!\!\!\!\!=\frac{\sqrt{p_{n}a_{nm}}\theta^{*}\mathbb{R}_{nm}}{\sum_{j\in\mathcal{C}_{n}}{p_{j}a_{jm}}\mathbb{R}_{jm}+\sigma^{2}\mathbb{I}_{L}}\mathbb{y}_{nm}^{\textrm{p}}$
@@ III-C Downlink Transmission
$\textbf{x}_{m}=\sum_{n=1}^{N}\sqrt{\varkappa_{nm}a_{mn}}\textbf{v}_{nm}s_{n}$ (15)
@@
$y_{n}$ $=$ $\sum_{m=1}^{M}\textbf{h}_{nm}^{H}\textbf{x}_{nm}s_{n}$ (18)
$\!\!\!\!\!\!\!\!\!\!=\sum_{m=1}^{M}\sqrt{\varkappa_{nm}a_{nm}}\textbf{h}_{nm}^{H}\textbf{v}_{nm}s_{n}$
$+\sum_{m=1}^{M}\sum_{\begin{subarray}{c}i=1\\ i\neq n\end{subarray}}^{N}\sqrt{\varkappa_{im}a_{im}}\textbf{h}_{im}^{H}\textbf{v}_{im}s_{i}+w_{n}$
@@
$\textbf{v}_{nm}=\theta_{nm}\hat{\mathbb{h}}_{nm}$ (19)
@@
$\textrm{SINR}_{n}=\frac{\Big{|}\mathbb{E}\Big{\{}\sum_{m=1}^{M}\sqrt{\varkappa_{nm}a_{nm}}\textbf{h}_{nm}^{H}\textbf{v}_{nm}\Big{\}}\Big{|}^{2}}{\sum_{m=1}^{M}{\sum_{\begin{subarray}{c}i=1\\ i\neq n\end{subarray}}^{N}\varkappa_{im}a_{im}}\mathbb{E}\Big{\{}\Big{|}\textbf{h}_{im}^{H}\textbf{v}_{im}\Big{|}^{2}\Big{\}}+w_{n}}$ (22)
Thus, the achievable spectral efficiency for the $n$-th user is given by
@@
×$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\times\log_{2}\big{(}1+\frac{\Big{|}\mathbb{E}\Big{\{}\sum_{m=1}^{M}\sqrt{\varkappa_{nm}a_{nm}}\theta_{nm}\textbf{h}_{nm}^{H}\hat{\mathbb{h}}_{nm}\Big{\}}\Big{|}^{2}}{E_{n}+F_{n}+w_{n}}\big{)}$
@@
$E_{n}=\textrm{var}\Big{\{}\sum_{m=1}^{M}\sqrt{\varkappa_{nm}a_{nm}}\theta_{nm}\textbf{h}_{nm}^{H}\hat{\mathbb{h}}_{nm}\Big{\}}$ (24)
$F_{n}=\sum_{m=1}^{M}{\sum_{\begin{subarray}{c}i=1\\ i\neq n\end{subarray}}^{N}\varkappa_{im}a_{im}}\theta_{nm}\mathbb{E}\Big{\{}\Big{|}\textbf{h}_{im}^{H}\hat{\mathbb{h}}_{im}\Big{|}^{2}\Big{\}}$ (27)
@@ IV Simulation Results
In Figs. 1(a) and 1(b), we plot the cumulative distribution function (CDF) of the per-user downlink spectral efficiency of the proposed UC-DMIM system for $M=100$ and $M=400$, respectively. These figures illustrate the system performance when the RSAP is selected based on the best channel criteria. As depicted in the figure, since the propagation delay phase can be perfectly estimated and incorporated into the precoding process, the system performance is significantly enhanced when this phase shift is compensated. For example, as shown in Fig. 1(b), the performance of the synchronous transmission with $L=2$ is almost the same as that of asynchronous transmission with $L=4$. This indicates that phase compensation can either enhance the system performance or reduce the number of distributed or co-located antenna requirements. Furthermore, it is clear from Figs. 1(a) and 1(b) that the system performance in terms of spectral efficiency is enhanced with increasing the number of orbiting satellites since this increases the probability of more satellites joining the user’s SC.
@@ V Conclusion
In this paper, we proposed a scalable distributed massive MIMO system for LEO satellite networks using dynamic user-centric clustering. The proposed UC-DMIMO system underscores its effectiveness in improving the downlink spectral efficiency and handover rate. Notably, compensating for propagation delay phase shifts significantly enhances system efficiency, either by boosting the downlink data rate or reducing antenna requirements. Each user is assigned an RSAP which shares the user information with the other satellites in its SC. Two distinct methods are introduced for RSAP selection; either opting for the LEO satellite with the best channel or choosing the one with the maximum service time. The comparison between RSAP selection methods further reinforces the system’s superiority over a baseline NCT systems. Moreover, with almost the same spectral efficiency performance as the FC-DMIMO system, the UC-DMIMO provides significantly smaller cluster size which reduces the backhaul overhead requirements. Overall, these findings affirm the potential of the proposed UC-DMIMO architecture to elevate system performance in terms of spectral efficiency, handover rate, and scalable implementation in cooperative LEO satellite networks.

### v2

@@ Abstract
LEO satellite networks, distributed mMIMO, user-centric, dynamic clustering, asynchronous transmission.
@@ I Introduction
Multiple-input multiple-output (MIMO) is a technology implemented to enhance the performance in wireless communication systems [1]. However, co-located MIMO in satellites communication fails to provide the same performance as that in terrestrial networks [2]. This is because satellite channels generally differ from those of the terrestrial network. In particular, strong line-of-sight (LoS) channel is one of the main characteristics of satellite channels due to the lack of scatters in the space. This satellite channel characteristic results in high spatial correlation when co-located MIMO systems are used, making it difficult for the receiver to estimate the channel coefficient for each antenna and hence affecting the validity of the channel state information which is necessary for MIMO system operation. As such, satellite networks, have the drawback in that they are unable to effectively exploit the performance enhancement provided by MIMO technology compared to terrestrial networks [3]. To overcome this issue, instead of using co-located MIMO, a distributed MIMO system is proposed which is able to provide uncorrelated channels among spatially separated antennas [4].
Distributed MIMO in space requires the cooperation of satellites to enable beamforming or joint transmission and reception. In [5], a joint transmission scheme is investigated using two LEO satellites. It is shown that the relative positioning of the cooperating satellites and the ground user substantially impacts system performance due to the predominance of LOS channel. The authors in [6] introduced orthogonal time frequency space (OTFS) for distributed multi-satellite cooperative transmission scheme to effectively enhance downlink performance. The problem of overlap in satellite coverage areas is studied in [7] where a user receives multiple signals in a multi-satellite scenario. In [4] a distributed massive MIMO (mMIMO) system is proposed for LEO satellite networks where multiple satellite access points cooperate to serve a set of ground users. A theoretical framework to investigate the spectral efficiency performance of distributed mMIMO in space was developed in [8]. The results in [4] and [8] show that in addition to improving the average system spectral efficiency, distributed mMIMO can reduce the handover rate since the user receives its data from a cluster of serving satellites instead of a single satellite connection. In [9], a distributed beamforming technique was proposed to support direct communications from LEO satellites to smartphones. To enhance the user quality of service (QoS), a coordinated multi-satellite joint transmission is investigated in [10]. The study demonstrates that this system not only achieves higher coverage probability than traditional single-satellite systems but also optimizes user’s ergodic rates.
@@
In this paper, we propose a scalable distributed mMIMO system designed for LEO satellite networks, utilizing dynamic user-centric clustering. The proposed system is called a user-centric distributed mMIMO (UC-DMIMO) system. With a focus on achieving scalable implementation, we provide new algorithms for initial cooperative access, pilot assignment, and cluster selection. To further enhance system efficiency in terms of complexity and handover rate, each user is allocated a reference satellite access point (RSAP) responsible for exchanging the user’s information with other satellites in its designated serving cluster (SC). Furthermore, phase-aware pre-coding is used to synchronize the phases of signals received from different spatially distributed antennas. Finally, the system performance in terms of spectral efficiency and the SC size of the proposed UC-DMIMO system is compared with that of the baseline full-cooperative DMIMO (FC-DMIMO) systems, where all satellites cooperatively serve all users simultaneously, as well as non-cooperative transmission (NCT) systems, where each user is connected to only a single satellite.
@@ II-B Channel Model
$\textbf{h}_{nm}=\beta_{nm}\textbf{g}_{nm},$ (1)
@@
where $\kappa$ refers to the Rician K-factor and $\textbf{g}^{\prime}_{nm}$ and $\textbf{g}^{\prime\prime}_{nm}$ corresponds to the LOS and NLOS components. Specifically, $\textbf{g}^{\prime\prime}_{nm}$ characterizes the random component, accounting for the scattered multipath that follows a Rayleigh distribution its elements are independent and identically distributed random variables with a zero-mean and unit variance. Besides, according to [12], the $l$-th element of $\textbf{g}^{\prime}_{nm}$ is given as
@@
$\beta_{nm}=\frac{1}{{\mathcal{L}^{prop}_{nm}}\mathcal{L}^{sh}_{nm}\mathcal{L}^{ant}_{nm}\mathcal{L}^{other}_{nm}},$ (4)
@@
${\mathcal{L}^{prop}_{nm}}[\textrm{dB}]=32.45+20\textrm{log}_{10}f+20\textrm{log}_{10}r_{nm}.$ (5)
The parameters $\mathcal{L}^{sh}_{nm}$ in (4) represents the losses due to the shadowing effect, modeled as a log-normal random variable, i.e., $\mathcal{L}^{sh}_{nm}\sim\mathcal{N}(0,\sigma_{sh}^{2})$, where $\sigma_{sh}^{2}$ is the shadowing variance. $\mathcal{L}^{ant}_{ml}$ denotes the gain losses due to beam misalignment . Based on the 3GPP report in [11], this losses is modeled as
$\mathcal{L}^{ant}_{nm}=\frac{1}{4}\bigg{|}\frac{2\pi\eta\sin{w_{nm}}}{J_{1}(2\pi\eta\sin{w_{nm}})}\bigg{|}^{2},$ (6)
@@ II-C Asynchronous Reception
$\theta_{nm}=\exp{\big{(}-2\pi\frac{\Delta t_{nm}}{T_{s}}\big{)}},$ (7)
@@ III-B Channel Estimation
$\mathbb{Y}^{P}_{m}=\sum_{n=1}^{N}\sqrt{a_{nm}p_{n}}\tilde{\mathbb{h}}_{nm}\mathbb{\Xi}_{n}^{T}+\mathbb{W}_{m},$ (9)
@@
$\!\!\!\!\!\!\!\!\!\!=\frac{\sqrt{p_{n}a_{nm}}\theta^{*}\mathbb{R}_{nm}}{\sum_{j\in\mathcal{C}_{n}}{p_{j}a_{jm}}\mathbb{R}_{jm}+\sigma^{2}\mathbb{I}_{L}}\mathbb{y}_{nm}^{\textrm{p}},$
@@ III-C Downlink Transmission
$\textbf{x}_{m}=\sum_{n=1}^{N}\sqrt{\varkappa_{nm}a_{mn}}\textbf{v}_{nm}s_{n},$ (15)
@@
$y_{n}$ $=$ $\sum_{m=1}^{M}\textbf{h}_{nm}^{H}\textbf{x}_{nm}s_{n}=\sum_{m=1}^{M}\sqrt{\varkappa_{nm}a_{nm}}\textbf{h}_{nm}^{H}\textbf{v}_{nm}s_{n}$ (16)
$+\sum_{m=1}^{M}\sum_{\begin{subarray}{c}i=1\\ i\neq n\end{subarray}}^{N}\sqrt{\varkappa_{im}a_{im}}\textbf{h}_{im}^{H}\textbf{v}_{im}s_{i}+w_{n},$
@@
$\textbf{v}_{nm}=\theta_{nm}\hat{\mathbb{h}}_{nm}.$ (17)
@@
$\textrm{SINR}_{n}=\frac{\Big{|}\mathbb{E}\Big{\{}\sum_{m=1}^{M}\sqrt{\varkappa_{nm}a_{nm}}\textbf{h}_{nm}^{H}\textbf{v}_{nm}\Big{\}}\Big{|}^{2}}{\sum_{m=1}^{M}{\sum_{\begin{subarray}{c}i=1\\ i\neq n\end{subarray}}^{N}\varkappa_{im}a_{im}}\mathbb{E}\Big{\{}\Big{|}\textbf{h}_{im}^{H}\textbf{v}_{im}\Big{|}^{2}\Big{\}}+w_{n}}.$ (18)
Thus, the achievable spectral efficiency for the $n$-th user is
@@
×$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\times\log_{2}\big{(}1+\frac{\Big{|}\mathbb{E}\Big{\{}\sum_{m=1}^{M}\sqrt{\varkappa_{nm}a_{nm}}\theta_{nm}\textbf{h}_{nm}^{H}\hat{\mathbb{h}}_{nm}\Big{\}}\Big{|}^{2}}{E_{n}+F_{n}+w_{n}}\big{)},$
@@
$E_{n}=\textrm{var}\Big{\{}\sum_{m=1}^{M}\sqrt{\varkappa_{nm}a_{nm}}\theta_{nm}\textbf{h}_{nm}^{H}\hat{\mathbb{h}}_{nm}\Big{\}},$ (20)
$F_{n}=\sum_{m=1}^{M}{\sum_{\begin{subarray}{c}i=1\\ i\neq n\end{subarray}}^{N}\varkappa_{im}a_{im}}\theta_{nm}\mathbb{E}\Big{\{}\Big{|}\textbf{h}_{im}^{H}\hat{\mathbb{h}}_{im}\Big{|}^{2}\Big{\}}.$ (21)
@@ IV Simulation Results
In Figs. 2(a) and 2(b), we plot the cumulative distribution function (CDF) of the per-user downlink spectral efficiency of the proposed UC-DMIM system for $M=100$ and $M=400$, respectively. These figures illustrate the system performance when the RSAP is selected based on the best channel criteria. As depicted in the figure, since the propagation delay phase can be perfectly estimated and incorporated into the precoding process, the system performance is significantly enhanced when this phase shift is compensated. For example, as shown in Fig. 2(b), the performance of the synchronous transmission with $L=2$ is almost the same as that of asynchronous transmission with $L=4$. This indicates that phase compensation can either enhance the system performance or reduce the number of distributed or co-located antenna requirements. Furthermore, it is clear from Figs. 2(a) and 2(b) that the system performance in terms of spectral efficiency is enhanced with increasing the number of orbiting satellites since this increases the probability of more satellites joining the user’s SC.
@@ V Conclusion
In this paper, we proposed a scalable distributed massive MIMO system for LEO satellite networks using dynamic user-centric clustering. The proposed UC-DMIMO system underscores its effectiveness in improving the downlink spectral efficiency and handover rate. Notably, compensating for propagation delay phase shifts significantly enhances system efficiency, either by boosting the downlink data rate or reducing antenna requirements. Each user is assigned an RSAP which shares the user information with the other satellites in its SC. Two distinct methods are introduced for RSAP selection; either opting for the LEO satellite with the best channel or choosing the one with the maximum service time. The comparison between RSAP selection methods further reinforces the system’s superiority over a baseline NCT systems. Moreover, with almost the same spectral efficiency performance as the FC-DMIMO system, the UC-DMIMO provides significantly smaller cluster size which reduces the backhaul overhead requirements. Overall, these findings affirm the potential of the proposed UC-DMIMO architecture to elevate system performance in terms of spectral efficiency, handover rate, and scalable implementation.